OpenStudy (anonymous):

. Solve for x x = 19 x = 12 x = 9 x = 16

OpenStudy (anonymous):

Please help!!

OpenStudy (cwrw238):

the 2 triangles ABC and DEC are similar so corresponding sides are in the same ratio so 7 / 3 = 12+(x - 3) / 12 7/3 = x + 9 / 12 84 = 3x + 27 can u finish this off - what is value of x?

OpenStudy (anonymous):

Ok so 57/3 = 19 or X=19 :) @cwrw238

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thanks :)